if a and b are complementary angles and a is 10 degrees greater than b, what size is angle b?
Answer by drk(1908)

We know that
a + b = 90
a = b + 10
By substitution we get
b+ 10 + b = 90
2b + 10 = 90
2b = 80
b = 40.